OpenStudy (bbb911):

The direction for mixing an insect spray say to uses 2 3/4 ounces of chemicals in each gallon of water. How many ounces of chemical should be mixed with 6 3/5 gallons of water? First, estimate the answer to the application problem then find the exact answer.

jimthompson5910 (jim_thompson5910):

2 3/4 = 2+3/4 2 3/4 = 8/4+3/4 2 3/4 = 11/4

jimthompson5910 (jim_thompson5910):

6 3/5 = 6+3/5 6 3/5 = 30/5+3/5 6 3/5 = 33/5

jimthompson5910 (jim_thompson5910):

you'll then set up the following proportion (amount of chemicals)/(amount of water) = (amount of chemicals)/(amount of water) (11/4)/(1) = (x)/(33/5) then cross multiply to get (11/4)*(33/5)= 1*x keep going to find x

OpenStudy (bbb911):

okay the estimated answer 18 right ?

jimthompson5910 (jim_thompson5910):

yes, very good

jimthompson5910 (jim_thompson5910):

the exact is what?

OpenStudy (bbb911):

this might take a moment give me a sec

jimthompson5910 (jim_thompson5910):

alright

OpenStudy (bbb911):

okay is it 18 3/20?

jimthompson5910 (jim_thompson5910):

you nailed it

OpenStudy (bbb911):

omgggggg!! yaay!!! :D

jimthompson5910 (jim_thompson5910):

lol

OpenStudy (bbb911):

thank you sooo much!! :)

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (bbb911):

u r a hero

jimthompson5910 (jim_thompson5910):

thanks, I appreciate the compliment

OpenStudy (bbb911):

welcome :)